OpenStudy (anonymous):

Solve by using matrices. If no solution, write inconsistent. (Just trying to double check my work, don't need anyone to show their work) {1 1 1 2 2 -1 -1 -5 1 2 2 5}

OpenStudy (anonymous):

Assuming the matrix represents a system of linear equations: \[x_{1}=x _{4}\] \[x _{2}=-x _{3}-3x _{4}\]

OpenStudy (anonymous):

Do you know what x, y, and z equal? It looks like z is arbitrary. And yes, it is system of linear equations.

OpenStudy (anonymous):

wait, does this represent x+y+z=2, or x+y+z+2a=0?

OpenStudy (anonymous):

x+y+z=2

OpenStudy (anonymous):

oh, forget about my first anser then.

OpenStudy (anonymous):

x=-1, y+z=3

OpenStudy (anonymous):

Awesome! That's what I got too. Thanks